An economist is interested in examining how an individual’s cigarette consumption (C) may be influenced by the price for a pack of cigarettes (P) and the individual’s annual income (I). Using data from 50 individuals, she estimates a log-log model and obtains the following regression results.

ln(C)ˆln(C)^ = 3.90 1.25ln(P) + 0.18ln(I)
p-values = (0.000) (0.005) (0.400)

a. Interpret the value of the elasticity of demand for cigarettes with respect to price. (Round your answer to 2 decimal places.)


b. At the 5% significance level, is the price elasticity of demand statistically significant?


  • Yes, since the p-value is less than the significance level.

  • Yes, since the p-value is not less than the significance level.

  • No, since the p-value is less than the significance level.

  • No, since the p-value is not less than the significance level.


c. Interpret the value of the income elasticity of demand for cigarettes. (Round your answer to 2 decimal places.)


d. At the 5% significance level, is the income elasticity of demand statistically significant?
Answer #1

a)

As the model shows, ln(C)^ln(C)^= 3.90-1.25ln(P) + 0.18ln(I)

This model shows the coefficient of price is -1.25.

Which means that for each unit the price of the pack cigarettes increases the consumption would decline by the 1.25%.

b)

option (a) is correct.

Reason: Since the p-value corresponding the price is 0.005, this shows that p-value<alpha-value(0.05)

And, when p-value is less than alpha value(level of significance), this means that we have sufficient evidence that the results are statistically significant.

c)

As the model shows, ln(C)^ln(C)^= 3.90-1.25ln(P) + 0.18ln(I)

This model shows the coefficient of price is 0.18.

This interprates that with increase in one unit of income would lead to 0.18% increase in the consumption of cigarettes.

d)

Since the p-value corresponding the price is 0.400, this shows that p-value>alpha-value(0.05)

And, when p-value is greater than alpha value(level of significance), this means that we do not have sufficient evidence that the results are statistically significant.

Income elasticity of demand is not statistically significant.